# Thermodynamic and observational implications of black holes in toroidal geometry

### v2

@@ Abstract
We investigate the thermodynamic and observational implications for the charged torus-like black holes, a class of solutions distinct from the classical Schwarzschild black holes. We explicitly derive the fundamental thermodynamic properties, such as heat capacity, P-V diagram, isothermal compressibility, Helmholtz free energy, and Gibbs free energy, under different entropy models. We find that only the exponential corrected entropy demonstrates multiple phase transitions, which we validate with the Ricci Scalar divergence obtained from the Ruppeiner formalism. This indicates that exponential corrected entropy is more sensitive to BH’s microstructure as compared to the Hawking-Bekenstein and Rènyi entropy models within the non-spherical (toroidal) horizon. In addition, we study the sparsity and emission rates of Hawking radiation, demonstrating that exponential correction entropy yields more consistent and stable behavior. In our observational analysis, we graphically demonstrate the behavior of redshift, blueshift, and gravitational shift, and identify specific conditions where the photon sphere radius exceeds the innermost stable circular orbit radius, which depends on the values of parameters such as electric charge and cosmological constant. The novel insight of this work is that despite this violation, our computed redshift, blueshift, and gravitational shifts fall within the range of the observational data of NGC 4258 and UGC 3789.
@@ I Introduction
Black holes (BHs) have become one of the most important topics in the theory of general relativity, which provides a very unique framework where the concepts of thermodynamics, gravity, and observational astrophysics align. Hawking and Bekenstein’s foundational work on BH thermodynamics has established a significant link between gravitational theory and quantum mechanics Hawking:1974rv ; Bekenstein:1973ur ; Bekenstein:1974ax ; Hawking:1975vcx ; Bekenstein:1972tm ; Bardeen:1973gs . In this framework, BHs follows the laws of thermodynamics, where surface gravity corresponds to temperature (Hawking temperature), the area of the event horizon signifies entropy and exhibiting phase transition akin to those in conventional thermodynamic system Bardeen:1973gs . Analytical frameworks like PVP-V criticality Chamblin:1999tk ; Chamblin:1999hg ; Kubiznak:2012wp ; Kubiznak:2016qmn ; Karch:2015rpa ; Mancilla:2024spp , stability analysis utilizing heat capacity, Gibbs free energy (GFE) and Helmholtz free energy (HFE) Rodrigues:2022qdp ; Simovic:2023yuv ; Ali:2019myr , provide significant insights on BH characteristics and their thermodynamic behavior. For example, a BH is considered locally and globally stable if the heat capacity, HFE, and GFE demonstrate positive behavior and vice versa (for further details regarding local and global stability of BHs, check Refs. Rodrigues:2022zph ; Dehyadegari:2017flm ; Wei:2020poh ; Liu:2014gvf ; Dehyadegari:2020tau ).
@@
In this context, thermodynamic geometry provides an effective framework to study the BH phase transition by introducing various thermodynamic metrics, and these metrics are constructed in terms of entropy, which gives us valuable insight regarding the phase transition when the divergence of Ricci curvature scalar determined from this metric formalism aligns with the zero point (ZP) or divergence of heat capacity. The metric formalism introduced by the Weinhold Weinhold:1975xej , based on equilibrium state space, which was later extended by the Ruppiener Ruppeiner:2008kd ; Ruppeiner:1995zz , whose formalism is closely linked with the Weinhold’s formalism (for other thermodynamic geometry formalism and their constructions check Refs. NaveenaKumara:2020biu ; Quevedo:2008ry ; Akbar:2011qw ; Hendi:2015cka ; Sarkar:2006tg ; Bhattacharya:2019qxe ; Hendi:2015xya ; Banerjee:2016nse ), it reveals valuable insights regarding the microscopic interaction and phase transitions. Although research on non-spherical BH configurations has increased, toroidal or torus-like BHs have not been investigated as extensively as spherically symmetric ones. The main reason is that they have very intricate topology and horizon geometry, which makes them difficult to analyze. These geometries naturally arise within AdS spacetime and offer a framework for studying the thermodynamic and observational characteristics. Departing from previous analyses, which focus on the hyperbolic and planar horizons, our toroidal framework features a novel dual periodic topology, providing modified thermodynamic curvature characteristics. In torus-like spacetimes, entropy correction plays a significant role because the traditional entropy (Hawking-Bekenstein entropy) formulated under classical assumptions may not fully incorporate quantum or topological complexities. To better describe these systems, Rènyi Odintsov:2023qfj and exponential corrected entropy ghosht introduce generic statistical or quantum modifications that may be fundamentally significant.
@@
This paper is arranged as follows: in Sec. II, we discuss the torus-like BH solution in detail. We analytically probe the thermodynamic aspects of charged torus-like BHs in Sec. III, where we derive thermodynamic quantities such as mass, temperature, and volume, analyze the heat capacity, PVP-V diagram, isothermal compressibility, GFE, and HFE in terms of various entropy models including Hawking-Bekenstein, Rènyi, and exponential corrected entropy frameworks. In Sec. IV, we utilize the thermodynamic geometry formalism to investigate the behavior of Ricci curvature scalar for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies to identify the possible phase transition. Sec. V provides a detailed analysis of the sparsity of the Hawking radiation and energy emission from charged torus-like BHs using the Hawking-Bekenstein, exponential corrected, and Rènyi entropies. Extending our thermodynamic analysis of charged torus-like BHs to the observational analysis, we analytically derive the behavior of a particle in time-like and null geodesics and employ these results to evaluate and graphically represent the frequency shift, including redshift, blueshift, and gravitational shift, in Sec. VI. The conclusion of our work is presented in Sec. VII.
@@ III.1 Thermodynamic Mass
Here, we determine the thermodynamic mass in terms of these entropy formalisms for the charged torus-like BH [^6] We want to mention here that the thermodynamics of charged torus-like BHs in terms of Hawking-Bekenstein entropy have been discussed in Refs. Ali:2023pyv ; Feng:2021vey ; Liang:2021elg ; Ditta:2023gin , but our aim is to discuss the impact of entropy corrections in comparison with the traditional entropy (Hawking-Bekenstein entropy). Therefore, we have analytically obtained the thermodynamic quantities in terms of the Hawking-Bekenstein, exponential correct, and Rènyi entropy models to enhance our understanding regarding the thermodynamic behavior of charged torus-like BHs.. Firstly, we compute the mass of the BH in terms of the Hawking-Bekenstein entropy formalism by using Eqs. (6) and (7), which yields
@@
Similarly, in the case of Rènyi entropy, one can easily derive the mass of the BH by employing Eqs. (6) and (10). So, the expression of mass in terms of the Rènyi entropy can be written as
@@
Here, we emphasize that $M(S_{\mathrm{Exp}},~P,~Q)$ and $M(S_{\mathrm{R}},~P,~Q)$ are not computed by directly plugging $S_{\mathrm{Exp}}$ and $S_{\mathrm{R}}$ into Eq. (6) (the Hawking-Bekenstein formula); instead, their derivation aligns with fundamental thermodynamic laws. We initiate the analysis of each entropy model by constructing their expressions as functions of the BH horizon radius $r_{\mathrm{e}}$; exponential corrected entropy is developed using the methodology provided in Ref. ghosht , and the Rènyi entropy is formulated via a non-extensive and non-additive approach from Ref. Odintsov:2023qfj . These expressions are subsequently rearranged to solve for $r_{\mathrm{e}}$ in terms of $S_{\mathrm{Exp}}$ or $S_{\mathrm{R}}$, respectively. By incorporating the inverted expressions introduced earlier into the torus-like BH’s metric function $f(r)$, the mass is re-expressed as a function of the corresponding entropy model. Thereby, $M_{\mathrm{Exp}}$ depends on the $S_{\mathrm{Exp}}$ and $M_{\mathrm{R}}$ depends on the $S_{\mathrm{R}}$. The approach complies with the first law of BH thermodynamics, thereby maintaining the physical validity of the derived results based on the chosen entropy model.
@@ III.2 Temperature
Before detailing the formulas for pressure and temperature within the exponential and Rènyi entropy approaches, the thermodynamic principles that underpin these derivations need to be properly explained. Under the generalized thermodynamic approach, different entropy formulations generate distinct thermodynamic phase spaces, compelling the formulation of BH mass within the context of each specific entropy model. As a result, the first law of BH thermodynamics in terms of exponential entropy can be altered accordingly, which is given as
@@
Similarly, the first law of BH thermodynamics is modified in terms of Rènyi entropy, which is given as
@@
We compute the thermodynamic temperature, volume, and potential associated with the Rènyi entropy for the charged torus-like BH, which are expressed as
@@
The foundation of the above-mentioned formalism lies in modern gravitational thermodynamics, which employs generalized entropic schemes such as the exponential corrected entropy and Rènyi entropy, capturing non-extensive and quantum gravitational features, respectively Rani:2025esb . We mention here that to determine $T_{\mathrm{Exp}}$ and $T_{\mathrm{R}}$, we employ a robust method involving the differentiation of the BH mass with respect to the relevant entropy functions. This ensures that our thermodynamic model is physically consistent and aligns with the foundational principles of generalized statistical mechanics.
@@ III.3 Local Stability
Similarly, we compute the heat capacity associated with the Rènyi entropy by utilizing Eq. (18) as
@@
Figure 1: $C_{P}$ versus the Hawking-Bekenstein, exponential corrected and Rényi entropy $S$ by setting $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We obtain various curves for different entropy models to notice the impact of these entropies on $C_{P}$, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are demonstrated as red solid, blue dotted, and black dotted-dashed trajectories, respectively.
Analyzing how specific heat varies with the size of a BH offers key insights into its stability under thermodynamic conditions and possible phase transitions. When the heat capacity is positive, the system tends to be stable and capable of equilibrating thermally with its surroundings; a negative value, on the other hand, points to instability, and a value of zero corresponds to a critical phase transition point. The heat capacity behavior in terms of different entropy frameworks is presented in Fig. 1 by setting $P=1,~\alpha=1$, and $Q=1$. Notably, we obtain different trajectories with respect to the entropy models that we employed in our analysis, such as for the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve), and Rènyi entropy (black dotted-dashed curve). We want to mention here that in the legends of all the thermodynamic figures, we present the Hawking-Bekenstein and exponential corrected entropies by HBE and Exp, respectively. It observes that heat capacity in terms of Hawking-Bekenstein and Rènyi entropies initially shows negative behavior for small values of entropy, and then they show positive behavior as the entropy rises, and also for small values, we obtain ZP. For example, there is only one ZP or phase transition in the cases of the Hawking-Bekenstein (red solid curve) and Rènyi entropy (black dotted-dashed curve), which are at $2.2214$ and $1.1698$, respectively. However, in the case of exponential entropy (blue dotted curve), the heat capacity initially decreases with a positive behavior, and for small intervals of entropy, it becomes negative; however, after that, it again begins to grow with a positive behavior. In addition, we obtain two ZPs at $1$ and $3.4674$, indicating that in the presence of exponential entropy, we can obtain more ZPs compared to other entropy frameworks. In BH thermodynamics, ZPs act as pivotal points that distinguish between two unique thermal states; for example, they separate the unstable region, in which heat capacity shows negative behavior, from the stable region, in which heat capacity presents positive behavior. It is also often referred to as the restriction point in BH thermodynamic analysis, beyond which the system’s behavior becomes stable. It is essential to mention here that by changing the parameter’s value location of ZP may also change; for example, in our case, if one chooses different values of Q,PQ,~P and $\alpha$ (in the case of Rènyi entropy), then the location of ZP might vary too.

### v3

@@ Abstract
We investigate the thermodynamic and observational implications for the charged torus-like black holes, a class of solutions distinct from the classical Schwarzschild black holes. We explicitly derive the fundamental thermodynamic properties, such as heat capacity, P-V diagram, isothermal compressibility, Helmholtz free energy, and Gibbs free energy, under different entropy models. We find that only the exponential corrected entropy demonstrates multiple phase transitions, which we validate with the Ricci Scalar divergence obtained from the Ruppeiner formalism. This indicates that exponential corrected entropy is more sensitive to BH’s microstructure as compared to the Hawking-Bekenstein and Rényi entropy models within the non-spherical (toroidal) horizon. In addition, we study the sparsity and emission rates of Hawking radiation, demonstrating that exponential correction entropy yields more consistent and stable behavior. In our observational analysis, we graphically demonstrate the behavior of redshift, blueshift, and gravitational shift, and identify specific conditions where the photon sphere radius exceeds the innermost stable circular orbit radius, which depends on the values of parameters such as electric charge and cosmological constant. The novel insight of this work is that despite this violation, our computed redshift, blueshift, and gravitational shifts fall within the range of the observational data of NGC 4258 and UGC 3789.
@@ I Introduction
Black holes (BHs) have become one of the most important topics in the theory of general relativity, which provides a very unique framework where the concepts of thermodynamics, gravity, and observational astrophysics align. Hawking and Bekenstein’s foundational work on BH thermodynamics has established a significant link between gravitational theory and quantum mechanics Hawking:1974rv ; Bekenstein:1973ur ; Bekenstein:1974ax ; Hawking:1975vcx ; Bekenstein:1972tm ; Bardeen:1973gs . In this framework, BHs follows the laws of thermodynamics, where surface gravity corresponds to temperature (Hawking temperature), the area of the event horizon signifies entropy and exhibiting phase transition akin to those in conventional thermodynamic system Bardeen:1973gs . Analytical frameworks like PVP-V criticality Chamblin:1999tk ; Chamblin:1999hg ; Kubiznak:2012wp ; Kubiznak:2016qmn ; Karch:2015rpa ; Mancilla:2024spp , stability analysis utilizing heat capacity, Gibbs free energy (GFE) and Helmholtz free energy (HFE) Rodrigues:2022qdp ; Simovic:2023yuv ; Ali:2019myr ; Rani:2025esb , provide significant insights on BH characteristics and their thermodynamic behavior. For example, a BH is considered locally and globally stable if the heat capacity, HFE, and GFE demonstrate positive behavior and vice versa (for further details regarding local and global stability of BHs, check Refs. Rodrigues:2022zph ; Dehyadegari:2017flm ; Wei:2020poh ; Liu:2014gvf ; Dehyadegari:2020tau ).
@@
In this context, thermodynamic geometry provides an effective framework to study the BH phase transition by introducing various thermodynamic metrics, and these metrics are constructed in terms of entropy, which gives us valuable insight regarding the phase transition when the divergence of Ricci curvature scalar determined from this metric formalism aligns with the zero point (ZP) or divergence of heat capacity. The metric formalism introduced by the Weinhold Weinhold:1975xej , based on equilibrium state space, which was later extended by the Ruppiener Ruppeiner:2008kd ; Ruppeiner:1995zz , whose formalism is closely linked with the Weinhold’s formalism (for other thermodynamic geometry formalism and their constructions check Refs. NaveenaKumara:2020biu ; Quevedo:2008ry ; Akbar:2011qw ; Hendi:2015cka ; Sarkar:2006tg ; Bhattacharya:2019qxe ; Hendi:2015xya ; Banerjee:2016nse ), it reveals valuable insights regarding the microscopic interaction and phase transitions. Although research on non-spherical BH configurations has increased, toroidal or torus-like BHs have not been investigated as extensively as spherically symmetric ones. The main reason is that they have very intricate topology and horizon geometry, which makes them difficult to analyze. These geometries naturally arise within AdS spacetime and offer a framework for studying the thermodynamic and observational characteristics. Departing from previous analyses, which focus on the hyperbolic and planar horizons, our toroidal framework features a novel dual periodic topology, providing modified thermodynamic curvature characteristics. In torus-like spacetimes, entropy correction plays a significant role because the traditional entropy (Hawking-Bekenstein entropy) formulated under classical assumptions may not fully incorporate quantum or topological complexities. To better describe these systems, Rényi Odintsov:2023qfj and exponential corrected entropy ghosht introduce generic statistical or quantum modifications that may be fundamentally significant.
@@
This paper is arranged as follows: in Sec. II, we discuss the torus-like BH solution in detail. We analytically probe the thermodynamic aspects of charged torus-like BHs in Sec. III, where we derive thermodynamic quantities such as mass, temperature, and volume, analyze the heat capacity, PVP-V diagram, isothermal compressibility, GFE, and HFE in terms of various entropy models including Hawking-Bekenstein, Rényi, and exponential corrected entropy frameworks. In Sec. IV, we utilize the thermodynamic geometry formalism to investigate the behavior of Ricci curvature scalar for the Hawking-Bekenstein, exponential corrected, and Rényi entropies to identify the possible phase transition. Sec. V provides a detailed analysis of the sparsity of the Hawking radiation and energy emission from charged torus-like BHs using the Hawking-Bekenstein, exponential corrected, and Rényi entropies. Extending our thermodynamic analysis of charged torus-like BHs to the observational analysis, we analytically derive the behavior of a particle in time-like and null geodesics and employ these results to evaluate and graphically represent the frequency shift, including redshift, blueshift, and gravitational shift, in Sec. VI. The conclusion of our work is presented in Sec. VII.
@@ III Thermodynamics Quantities of Torus-like BH Through the Corrected Entropies
It is important to mention that the thermodynamics models motivated by the quantum gravity indicate that the generalized entropy given in Refs. Nojiri:2022aof ; Nojiri:2022dkr ; Nojiri:2022sfd ; Elizalde:2025iku ; Odintsov:2023vpj ; Nojiri:2024zdu ; Odintsov:2022qnn provides the most comprehensive parametrization of horizon entropy, incorporating the logarithmic, power-law, and non-perturbative deviations from the Bekenstein-Hawking area law. The generality of this framework ensures that Rényi, Hawking-Bekenstein, Sharma-Mittal, Barrow, Tsallis, Loop quantum gravity, and Kindeski entropies arise with its specific limits. Despite offering significant mathematical stability, the generalized framework contains numerous free parameters that cannot be uniquely fixed, complicating the calculation of thermodynamic quantities and introducing ambiguity in thermodynamic studies. Therefore, using this generalized entropy framework makes our thermodynamic analysis more intricate, and it is also not analytically invertible.
For these reasons, in this paper, we primarily focus on these representative and analytically tractable entropy frameworks, which are indeed the subclasses of the generalized entropy given in Refs. Nojiri:2022aof ; Nojiri:2022dkr ; Nojiri:2022sfd ; Elizalde:2025iku ; Odintsov:2023vpj ; Nojiri:2024zdu ; Odintsov:2022qnn ; the Hawking-Bekenstein and Rényi entropies, which capture the non-extensive effects associated with horizons with non-trivial topology, and the exponential corrected entropy, which accounts for quantum gravity corrections. Each of these entropy expressions arises from a specific limit of the generalized entropy framework, while still maintaining the analytic mapping $S({r_{\mathrm{e}}})\leftrightarrow r_{\mathrm{e}}(S)$ (which is quite intricate in the case of generalized entropy), allowing us to derive the expression for mass, conjugate temperature with respect to the corresponding entropy, heat capacity, P-V diagram, isotherm compressibility, free energies (HFE and GFE), and Ruppeiner curvature. Also, since in our analysis we want to validate the divergence in heat capacity using the thermodynamic geometry metric formalisms, the generalized entropy complicates the derivation and validation. Thus, the choice of entropy models in our analysis maintains physical comprehensiveness while remaining numerically tractable, enabling us to make a clear and straightforward comparison of thermodynamic features corresponding with the Hawking-Bekenstein, Rényi, and exponential corrected entropies in the case of charged torus-like BHs.
@@ III.1 Thermodynamic Mass
Here, we determine the thermodynamic mass in terms of these entropy formalisms for the charged torus-like BH [^6] We want to mention here that the thermodynamics of charged torus-like BHs in terms of Hawking-Bekenstein entropy have been discussed in Refs. Ali:2023pyv ; Feng:2021vey ; Liang:2021elg ; Ditta:2023gin , but our aim is to discuss the impact of entropy corrections in comparison with the traditional entropy (Hawking-Bekenstein entropy). Therefore, we have analytically obtained the thermodynamic quantities in terms of the Hawking-Bekenstein, exponential correct, and Rényi entropy models to enhance our understanding regarding the thermodynamic behavior of charged torus-like BHs.. Firstly, we compute the mass of the BH in terms of the Hawking-Bekenstein entropy formalism by using Eqs. (6) and (7), which yields
@@
Similarly, in the case of Rényi entropy, one can easily derive the mass of the BH by employing Eqs. (6) and (10). So, the expression of mass in terms of the Rényi entropy can be written as
@@
Here, we emphasize that $M(S_{\mathrm{Exp}},~P,~Q)$ and $M(S_{\mathrm{R}},~P,~Q)$ are not computed by directly plugging $S_{\mathrm{Exp}}$ and $S_{\mathrm{R}}$ into Eq. (6) (the Hawking-Bekenstein formula); instead, their derivation aligns with fundamental thermodynamic laws. We initiate the analysis of each entropy model by constructing their expressions as functions of the BH horizon radius $r_{\mathrm{e}}$; exponential corrected entropy is developed using the methodology provided in Ref. ghosht , and the Rényi entropy is formulated via a non-extensive and non-additive approach from Ref. Odintsov:2023qfj . These expressions are subsequently rearranged to solve for $r_{\mathrm{e}}$ in terms of $S_{\mathrm{Exp}}$ or $S_{\mathrm{R}}$, respectively. By incorporating the inverted expressions introduced earlier into the torus-like BH’s metric function $f(r)$, the mass is re-expressed as a function of the corresponding entropy model. Thereby, $M_{\mathrm{Exp}}$ depends on the $S_{\mathrm{Exp}}$ and $M_{\mathrm{R}}$ depends on the $S_{\mathrm{R}}$. The approach complies with the first law of BH thermodynamics, thereby maintaining the physical validity of the derived results based on the chosen entropy model.
@@ III.2 Temperature
Before detailing the formulas for pressure and temperature within the exponential and Rényi entropy approaches, the thermodynamic principles that underpin these derivations need to be properly explained. Under the generalized thermodynamic approach, different entropy formulations generate distinct thermodynamic phase spaces, compelling the formulation of BH mass within the context of each specific entropy model. As a result, the first law of BH thermodynamics in terms of exponential entropy can be altered accordingly, which is given as
@@
Similarly, the first law of BH thermodynamics is modified in terms of Rényi entropy, which is given as
@@
We compute the thermodynamic temperature, volume, and potential associated with the Rényi entropy for the charged torus-like BH, which are expressed as
@@
The foundation of the above-mentioned formalism lies in modern gravitational thermodynamics, which employs generalized entropic schemes such as the exponential corrected entropy and Rényi entropy, capturing non-extensive and quantum gravitational features, respectively Rani:2025esb . We mention here that to determine $T_{\mathrm{Exp}}$ and $T_{\mathrm{R}}$, we employ a robust method involving the differentiation of the BH mass with respect to the relevant entropy functions. This ensures that our thermodynamic model is physically consistent and aligns with the foundational principles of generalized statistical mechanics.
Importantly, employing a generalized entropy does not imply any correction or modification to the Hawking temperature, which stays determined entirely by the surface gravity through the relation $T_{\mathrm{H}}=\kappa/2\pi$. As emphasized in Ref. Nojiri:2021czz , the inconsistency originates from substituting a non-Hawking-Bekenstein entropy into the first law without modifying the usual Bekenstein-Hawking mass-radius relation. By applying this procedure, the temperature becomes inaccurate since the function of mass is no longer conjugate to the modified entropy model within thermodynamics. In our approach, we avoid this inconsistency by computing the thermodynamic mass for each entropy model directly from the horizon condition and subsequently evaluating the temperature by using the first law formula given in Eqs. (15) and (17). Regardless of whether the entropy is the Hawking-Bekenstein, Rényi, or exponential corrected, we first reconstruct the $M(S)$ by utilizing the specific entropy-radius relation and then determine the conjugate temperature associated with the corresponding entropy by applying the relation given in Eqs. (14), (16), and (18). This guarantees that the temperature used in our thermodynamic treatment is consistent with the generalized entropy, resolving the inconsistency highlighted in Ref. Nojiri:2021czz , and this approach is adopted for different BHs as given in Refs . Zafar:2025sxl ; Nakarachinda:2022gsb ; Bhattacharjee:2025oje ; Yasir:2024zgm ; Yasir:2025uov ; Rani:2025esb ; Tariq:2025wiy ; Toledo:2019amt ; Bhattacharya:2024bjp ; Ladghami:2024sen .
@@ III.3 Local Stability
Similarly, we compute the heat capacity associated with the Rényi entropy by utilizing Eq. (18) as
@@
Figure 1: $C_{P}$ versus the Hawking-Bekenstein, exponential corrected and Rényi entropy $S$ by setting $P=1,~Q=1,~\alpha=1$ (in Rényi entropy case). We obtain various curves for different entropy models to notice the impact of these entropies on $C_{P}$, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rényi entropies are demonstrated as red solid, blue dotted, and black dotted-dashed trajectories, respectively.
Analyzing how specific heat varies with the size of a BH offers key insights into its stability under thermodynamic conditions and possible phase transitions. When the heat capacity is positive, the system tends to be stable and capable of equilibrating thermally with its surroundings; a negative value, on the other hand, points to instability, and a value of zero corresponds to a critical phase transition point. The heat capacity behavior in terms of different entropy frameworks is presented in Fig. 1 by setting $P=1,~\alpha=1$, and $Q=1$. Notably, we obtain different trajectories with respect to the entropy models that we employed in our analysis, such as for the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve), and Rényi entropy (black dotted-dashed curve). We want to mention here that in the legends of all the thermodynamic figures, we present the Hawking-Bekenstein and exponential corrected entropies by HBE and Exp, respectively. It observes that heat capacity in terms of Hawking-Bekenstein and Rényi entropies initially shows negative behavior for small values of entropy, and then they show positive behavior as the entropy rises, and also for small values, we obtain ZP. For example, there is only one ZP or phase transition in the cases of the Hawking-Bekenstein (red solid curve) and Rényi entropy (black dotted-dashed curve), which are at $2.2214$ and $1.1698$, respectively. However, in the case of exponential entropy (blue dotted curve), the heat capacity initially decreases with a positive behavior, and for small intervals of entropy, it becomes negative; however, after that, it again begins to grow with a positive behavior. In addition, we obtain two ZPs at $1$ and $3.4674$, indicating that in the presence of exponential entropy, we can obtain more ZPs compared to other entropy frameworks. In BH thermodynamics, ZPs act as pivotal points that distinguish between two unique thermal states; for example, they separate the unstable region, in which heat capacity shows negative behavior, from the stable region, in which heat capacity presents positive behavior. It is also often referred to as the restriction point in BH thermodynamic analysis, beyond which the system’s behavior becomes stable. It is essential to mention here that by changing the parameter’s value location of ZP may also change; for example, in our case, if one chooses different values of Q,PQ,~P and $\alpha$ (in the case of Rényi entropy), then the location of ZP might vary too.